A number is divided into two parts, such that one part is 12 more than the other. If the two parts are in the ratio 4 : 3, find the number and the two parts.
Answer by ikleyn(52786)

.
A number is divided into two parts, such that one part is 12 more than the other.
If the two parts are in the ratio 4 : 3, find the number and the two parts.
~~~~~~~~~~~~~~~~~~~~

From the problem, the parts are 4x and 3x, where x is some common measure, now unknown.


We also are given that

    4x - 3x = 12.


It gives us x = 12, so the parts are 4*12 = 48 and 3*12 = 36.


Thus the number is 48+36 = 84 and the parts are 48 and 36.    ANSWER

Solved.